Physics Shift Report¶
Dates: 04-10-2017 - 08-10-2017
Shifter: Christoper Hunt
Run Numbers: 09845 - 09904
Analysis Objectives¶
Run Number Table¶
Objective | Analysis Code | Channel | Setting | Run Numbers |
lH2 Scattering | S14a | Field Off | 3-170+M3-Test1 | 9854, 9856, 9859, 9861, 9882 |
S14b | Field Off | 3-200+M3-Test2 | 9855, 9857, 9858, 9860, 9881 | |
S14c | Field Off | 3-140+M3-Test3 | 9862, 9863, 9864, 9865, 9875, 9876, 9877, 9878, 9879, 9880 | |
lH2 Emittance Evolution | H30a | 2017-02-5 | 3-140+M3-Test3 | 9867 |
H30b | 2017-02-5 | 4-140+M3-Test2 | 9868, 9871, 9873, 9874 | |
H30c | 2017-02-5 | 10-140+M3-Test4 | 9869, 9870 | |
H31a | 2017-02-6 | 3-140+M3-Test3 | 9883, 9888, 9893, 9897, 9903 | |
H31b | 2017-02-6 | 4-140+M3-Test2 | 9884, 9885, 9889, 9894, 9898, 9904 | |
H31c | 2017-02-6 | 10-140+M3-Test4 | 9886, 9887, 9890, 9891, 9892, 9895, 9896, 9899, 9900, 9901, 9902 | |
Momentum Scan/Alignment | H29a | 2017-02-4 | 3-170+M3-Test1 | 9845, 9850 |
H29b | 2017-02-4 | 3-200+M3-Test2 | 9846, 9848, 9849, 9851 | |
H29c | 2017-02-4 | 3-120+M3-Test1 | 9847, 9852 |
Notes¶
Runs containing a field flip in SSDM2:
09867 - 09874
Elog, CDB, and Reconstruction¶
Elog entries: All ok.
CDB uploaded: All ok.
Online reconstruction: All ok.
Notes¶
M2D Flip¶
Friday 6th Oct.
Suspicious readings from the FC force metres, and a reduced transmission led the experts to believe that SSDM2 was powered in reverse.
Day was spent trying to validate that assumption using reconstructed tracks.
The residual between projected phi-angles and reconstructed phi-angles was used to exmine the data. In theory, if a coil is flipped, the particle track spends some time rotating in the opposite direction than it otherwise would. Therefore the reconstructed phi-angle would differ from the predicted. ( phi = atan2(py, px) )
Previous data sets showed residuals centred very close to zero - indicating that the number of rotations was approximately as expected.
Data from run 09867 showed a notable deviation from zero which was removed when a geometry containing a field flip was introduced. This was more than enough proof for the magnet experts to proceed in flipping the current leads to M2.
General¶
- There is a PMT power supply issue in the EMR plane 26. This is known about by the experts so no problem.
- lH2 Geometry within the reconstruction was fixed half way through the Physics Shifter cycle. (Runs 09895-09896) Take care with track extrapolation.
- Noticed that Pion Ref Run 09853 shows a suprisingly well focussed beam at TOF2. This is presumably a fluke of the optics+initial beam parameters?
Updated by Hunt, Christopher over 5 years ago · 11 revisions